# Enabling Natural Zero-Shot Prompting on Encoder Models via Statement-Tuning

### v2

@@ Abstract
While Large Language Models (LLMs) exhibit remarkable capabilities in zero-shot and few-shot scenarios, they often require computationally prohibitive sizes. Conversely, smaller Masked Language Models (MLMs) like BERT and RoBERTa achieve state-of-the-art results through fine-tuning but struggle with extending to few-shot and zero-shot settings due to their architectural constraints. Hence, we propose Statement-Tuning, a technique that models discriminative tasks as a set of finite statements and trains an Encoder model to discriminate between the potential statements to determine the label. We do Statement-Tuning on multiple tasks to enable cross-task generalization. Experimental results demonstrate that Statement-Tuning achieves competitive performance compared to state-of-the-art LLMs with significantly fewer parameters. Moreover, the study investigates the impact of several design choices on few-shot and zero-shot generalization, revealing that Statement-Tuning can achieve sufficient performance with modest training data and benefits from task and statement diversity for unseen task generalizability.
@@ 1 Introduction
Large Language Models (LLMs) have shown great capabilities in zero-shot and few-shot settings Radford et al. (2019); Brown et al. (2020); Artetxe et al. (2022). However, such capabilities are more difficult to observe in Encoder-only models like BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) due to their architectural design. These models are typically pretrained in an unsupervised manner on a large corpus with a Masked Language Modeling (MLM) Devlin et al. (2019) or Discriminative Clark et al. (2020) objective and fine-tuned by adding task-specific layers to enable their usage on a particular task such as binary/multi-label classification, token/sequence classification, multiple choice, etc. These task-specific layers, thus, can not be extended effectively to new tasks in a few-shot or zero-shot manner.
In this work, we explore the feasibility of utilizing encoder models that are usually specialized for a certain task to take on various, unseen Natural Language Understanding (NLU) tasks, akin to zero-shot prompting in decoder models. One benefit of using encoder models is that they are generally more compact. Yet, encoder models have achieved state-of-the-art results on many NLU tasks through task-specific finetuning, and thus it would be interesting if zero-shot/few-shot prompting could be adapted to be compatible with encoder models to leverage their powerful NLU capabilities at more computationally feasible sizes.
To address this issue, some techniques try to reformulate various downstream tasks with a unified format resembling the pretraining objective (MLM or Discriminative pretraining), enabling few-shot transfer for encoder models Xia et al. (2022); Schick and Schütze (2021a, b); Gao et al. (2021). These approaches tend to have multiple limitations including unintuitive/difficult implementations, complex handling of multi-token labels, and being unsuitable for zero-shot generalization.
In this work, we take inspiration from multitask instruction tuning methods for decoder models Wei et al. (2022a); Sanh et al. (2022) and unified format fine-tuning methods for encoder models to propose Statement-Tuning, a novel intuitive approach for encoder-only models to generalize to zero-shot and few-shot unseen tasks using universal multitask fine-tuning. Our approach thus has the generalization ability similar to decoder models.
@@
As seen in Figure 1, we verbalize a diverse set of NLU tasks into natural language statements, and then we fine-tune an encoder-only MLM, RoBERTa, on a universal binary sequence classification task, which we call Statement-Tuning, to assign a truth value (True or False) to any given statement. By fine-tuning encoder LLM across diverse tasks and statements, we show zero-shot generalization capabilities to unseen tasks by similarly transforming them into statements. Moreover, we show few-shot capabilities by continually fine-tuning this model with a small amount of downstream data, also formatted into statements. Statement-Tuning is capable of matching or even outperforming (32-shot and) zero-shot performance of many state-of-the-art LLMs with a fraction of the parameters.
Our ablation study shows that depending on the task, we can achieve sufficient few-shot and zero-shot generalizability with as few as 1,000 statements per training dataset or approximately 16,000 training statement examples in total, which correspond to even fewer original task examples since one example can be turned into multiple statements through different templates. Furthermore, we find that the proximity of the fine-tuning tasks to the evaluation tasks as well as the prompt and task diversity tend to have a beneficial effect on the effectiveness of Statement-Tuning.
@@
    To the best of our knowledge we are the first to enable natural, zero-shot task generalization in encoder models by verbalizing the input into statements and fine-tuning the model to perform binary classification on the truth value of a statement.
@@
    We expose that certain emergent abilities Wei et al. (2022b) like zero-shot generalization on unseen tasks previously thought to be exclusive to decoder-based LLMs can also be observed in much smaller encoder models when we do multitask Statement-Tuning.
@@
    We explore a large number of design choices to study how Statement-Tuning benefits from the number of statement examples and statement template and task diversity in multitask Statement-Tuning.
@@ 2 Related Work
Few-shot Approaches Using Encoder-Only Models
Prompt-based fine-tuning with cloze templates or label discrimination Schick and Schütze (2021a); Gao et al. (2021) effectively utilizes encoder models for few-shot learning. Prompts provide task context beyond traditional input-output training, enabling easier adaptation. However, multi-token labels and zero-shot generalization remain challenges Schick and Schütze (2021b); Tam et al. (2021). Our work addresses this by proposing a simpler reformulation method for improved unseen task generalization via multitask training.
Verbalizers in prompts leverage label semantics for task information (e.g., Tam et al. (2021)). Similar approaches include label conditioning tasks Tam et al. (2021) and TARS Halder et al. (2020). While TARS implicitly learns label-text relations, our method uses verbalization to explicitly connect text and labels in natural language statements. This facilitates generalization to unseen tasks without further fine-tuning, as the model learns a general semantic understanding rather than a rigid input structure.
Using Masked Language Models for Sequence Scoring
In a sense, Statement-Tuning can be viewed as fine-tuning MLMs for universal NLU sequence scoring based on the truth value of a statement and the subsequent generalization to any discriminative tasks. Previous approaches have explored using MLMs for general sequence scoring for non-autoregressive generation Wang and Cho (2019); Ghazvininejad et al. (2019). Salazar et al. (2020) use MLM pseudo-log-likelihood (PLL) scores as discriminators for generated output in Automatic Speech Recognition and low-resource Machine Translation and achieve improved generation through hypothesis re-scoring using the MLM PLL scores.
@@
LLMs excel at unseen-task/zero-shot generalization Brown et al. (2020). Building on this, recent work explores multitask training with diverse prompts for improved zero-shot performance Sanh et al. (2022); Wei et al. (2022a); Chung et al. (2022). These methods fine-tune large models on constructed datasets with various task prompts, achieving strong zero-shot results. However, effective instruction-tuned models often require billions of parameters Zhang et al. (2024), limiting their application to smaller models. Our work demonstrates achieving similar or superior generalization on smaller encoder-only models with less training data.
@@
#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion
Llama-2-7b-chat 7B 86.6 54.4 40.1 90.5 78.5 47.8 50.0
Mistral-7B-Instruct-v0.2 7B 89.4 73.0 41.4 88.9 82.3 57.7 55.3
Qwen1.5-7B-Chat 7B 87.0 75.5 42.1 95.3 79.7 59.1 57.8
Pythia-6.9B 6.9B 82.2 62.0 41.7 83.3 71.2 32.2 25.1
Pythia-2.8B 2.8B 79.6 68.4 41.2 77.7 69.7 12.1 35.4
Phi-2 2.7B 87.2 67.9 41.8 86.6 77.7 38.7 53.1
FlanT5-Large 770M 67.6 81.1 40.1 96.0 63.0 51.0 59.9
Qwen1.5-0.5B-Chat 500M 69.2 32.6 38.7 69.7 68.9 21.9 6.6
BART-large-mnli 406M 50.4 35.8 46.9 49.4 47.3 6.5 11.7
FlanT5-Small 60M 52.8 31.9 42.0 88.8 51.5 24.5 21.7
[…]
Our Approach: RoBERTa-base (4k) 125M 69.6 69.8 59.3 92.7 82.8 41.2 55.5
Our Approach: RoBERTa-large (Best) 355M 85.0 72.5 74.7 95.3 93.0 51.1 55.1
Our Approach: RoBERTa-large (50k) 355M 85.0 71.4 72.2 95.3 92.9 51.1 55.1
Full/3000-shot:
RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2
RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2
Table 1: Comparison of our approach against many pretrained open-source Encoder-Decoder and Decoder-only Pretrained Large Language Models on 7 Natural Language Understanding tasks in Zero-shot conditions. FT stands for Full Finetuning and is included for reference. We highlight all scores in red where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
@@ 3.1 Task Verbalization
Any discriminative task with a finite set of targets can be verbalized into a finite set of natural language statements. Figure 2 shows the example of converting the MNLI task into statements. Similar to prompting, each task has its own statement templates, based on each possible label. The truth label for training purposes on each statement depends on whether the statement contains the correct target label or not. We outline all the statement templates used for each dataset in Appendix A.
3.2 Statement Fine-Tuning
@@
To create the training data for statement fine-tuning, we exhaustively generate statements across 16 diverse NLP tasks using many varied statement templates per dataset: QQP Sharma et al. (2019), Winogrande Sakaguchi et al. (2020), PiQA Bisk et al. (2020), MNLI Williams et al. (2018), SNLI Bowman et al. (2015), Mintaka Sen et al. (2022), Yelp Polarity Zhang et al. (2015), WikiLingua Ladhak et al. (2020), SQuAD Rajpurkar et al. (2016), TweetEval’s Offensive task Zampieri et al. (2019), Massive FitzGerald et al. (2022); Bastianelli et al. (2020), Definite Pronoun Resolution Rahman and Ng (2012), QASC Khot et al. (2020), SciQ Johannes Welbl (2017), RACE Lai et al. (2017), and SAMSum Gliwa et al. (2019).
We explore different sample sizes for each dataset. We fine-tune RoBERTa Liu et al. (2019) with a binary sequence classification head to predict the truth value of the statements. By fine-tuning the model across diverse tasks, templates, and domains, the model should be able to generalize across unseen templates and tasks, as long as it can be phrased as a true/false statement.
3.3 Zero-Shot and Few-Shot Inference
@@
To perform inference on statement-finetuned RoBERTa, we also need to transform the input into statements. Specifically, we exhaustively generate a statement for each possible label, as shown in Figure 1. Then, for each statement corresponding to each label, we predict the probability of such a statement being true. The final label is the statement with the highest true probability. Zero-shot inference is done by directly performing the aforementioned inference regime on the statement-finetuned RoBERTa, while K-shot inference is done by first performing continual fine-tuning on K examples of task-specific statements.
@@ 4 Experimental Setup
4.1 Evaluation
In this work, we measure our model’s task generalizability using another set of 7 diverse datasets representing a variety of unseen tasks or unseen domains during training to judge the generalizability of our approach: Balanced COPA (BCOPA) Kavumba et al. (2019); Roemmele et al. (2011), MRPC Dolan and Brockett (2005), Emotion Saravia et al. (2018), Amazon Polarity McAuley and Leskovec (2013); Zhang et al. (2015), FigQA Liu et al. (2022), StoryCloze (2016) Yang et al. (2023), and Yahoo Answers Topics Zhang et al. (2015). Among the evaluation data, BCOPA, Emotion, FigQA, StoryCloze, and Yahoo Answers Topics are unseen tasks and hence examine the cross-task generalizability; however, MRPC and Amazon Polarity represent seen tasks but in different domains and demonstrate domain generalizability.
@@ 4.2 Statement Finetuning Configurations
We statement-finetune both RoBERTa-base and RoBERTa-large across diverse NLP tasks outlined in Section 3.2. However, as statement fine-tuning expands the dataset with various templates over all possible labels, it is arguably unwise to fine-tune on all possible generated statements. Moreover, each task has a different data size, leading to unbalanced fine-tuning data. Therefore, we sample statements randomly for each task, uniformly across true and false statements. We explore sample size from 1,000 statements to 50,000 statements for each dataset. Following best practices, we encourage the invariance to phrasing in diverse statements by designing multiple statement templates per dataset (a list of all statement templates is shown in Appendix A). We also explore the effect of statement diversity during training in Section 5.2.
After statement tuning is completed, we can further continue fine-tuning the model on the target downstream dataset. Specifically, we explore various n-shot configurations: Full/3,000-shot, 1,000-shot, 500-shot, 200-shot, and 32-shot, where we use limited data from the training sets of the corresponding dataset to fine-tune our statement-tuned models. For the Full/3,000-shot case, we cap the training set at 3,000 examples, otherwise, we use the entire set (this is the case for Amazon Polarity only). For StoryCloze, there is no training set, so we just carry out 32-shot (using 32 samples from the test set for fine-tuning and evaluating on the rest) and Zero-shot experiments. As for Yahoo Answers Topic and Emotion, due to them being multi-class classification tasks, we cap the n-shot analysis at 200-shot due to the larger number of choices per example (and hence a larger number of statements per example).

### v3

@@ Abstract
While Large Language Models (LLMs) exhibit remarkable capabilities in zero-shot and few-shot scenarios, they often require computationally prohibitive sizes. Conversely, smaller Masked Language Models (MLMs) like BERT and RoBERTa achieve state-of-the-art results through fine-tuning but struggle with extending to few-shot and zero-shot settings due to their architectural constraints. Hence, we propose Statement-Tuning, a technique that models discriminative tasks as a set of finite statements and trains an encoder model to discriminate between the potential statements to determine the label. We do Statement-Tuning on multiple tasks to enable cross-task generalization. Experimental results demonstrate that Statement-Tuning achieves competitive performance compared to state-of-the-art LLMs with significantly fewer parameters. Moreover, the study investigates the impact of several design choices on few-shot and zero-shot generalization, revealing that Statement-Tuning can achieve strong performance with modest training data and benefits from task and statement diversity for unseen task generalizability. We release all the code used to generate statement data, train and evaluate our Statement-Tuned models. [^1] Work In Progress
@@ 1 Introduction
Large Language Models (LLMs) have shown great capabilities in zero-shot and few-shot settings Radford et al. (2019); Brown et al. (2020); Artetxe et al. (2022). However, such capabilities are more difficult to observe in encoder-only models like BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) due to their architectural design. These models are typically pretrained in an unsupervised manner on a large corpus with a Masked Language Modeling Devlin et al. (2019) or Discriminative Clark et al. (2020) objective and fine-tuned by adding task-specific layers to enable their usage on a particular task such as binary/multi-label classification, token/sequence classification, multiple choice, etc. These task-specific layers, thus, can not be extended effectively to new tasks in a few-shot or zero-shot manner.
In this work, we explore the feasibility of utilizing encoder models that are usually specialized for a certain task to take on various, unseen Natural Language Understanding (NLU) tasks, akin to zero-shot prompting in decoder models. One benefit of using encoder models is that they are generally more compact. Yet, encoder models have achieved state-of-the-art results on many NLU tasks through task-specific fine-tuning. So it would be interesting if zero-shot/few-shot prompting could be adapted for encoder models to leverage their powerful NLU capabilities at more computationally feasible sizes.
To achieve this, some techniques try to reformulate various downstream tasks with a unified format resembling the pre-training objective, enabling few-shot transfer for encoder models Schick and Schütze (2021a, b); Xia et al. (2022). Without few-shot examples, the zero-shot generalization of these models relies mainly on the language modelling ability learned in the pre-training phase, not benefiting from further multitask training on diverse reformulated tasks.
In this work, we take inspiration from multitask instruction tuning methods for decoder models Wei et al. (2022a); Sanh et al. (2022) and unified format fine-tuning methods for encoder models Xu et al. (2023) to propose Statement-Tuning, a novel intuitive approach for encoder-only models to generalize to zero-shot and few-shot unseen tasks through universal multitask fine-tuning with data formatted as statements. Our approach thus has the generalization ability similar to decoder models with a fraction of the parameters and training data.
@@
As seen in Figure 1, we verbalize a diverse set of NLU tasks into natural language statements, and then fine-tune an encoder-only MLM, RoBERTa, on a universal binary sequence classification task, which we call Statement-Tuning, to assign a truth value (True or False) to any given statement. By fine-tuning encoder models across diverse tasks and statements, we show zero-shot generalization capabilities to unseen tasks by similarly transforming them into statements. Moreover, we show few-shot capabilities by continually fine-tuning this model with a small amount of downstream data, also formatted into statements. Statement-Tuning is capable of matching or even outperforming (32-shot and) zero-shot performance of many state-of-the-art LLMs with a fraction of the parameters.
Our ablation study shows that depending on the task, we can achieve substantial few-shot and zero-shot generalizability with as few as 1,000 statements per training dataset or approximately 16,000 training statement examples in total, which correspond to even fewer original task examples since one example can be turned into multiple statements through different templates. Furthermore, we find that the statement and task diversity tend to have a beneficial effect on the performance and generalizability of Statement-Tuning.
@@
    To the best of our knowledge, we are the first to propose a combination of elaborate statement formulation and Masked Language Models as a simple and effective data/resource-efficient alternative for LLMs for zero-shot NLU task generalization.
@@
    We expose that the emergent ability (Wei et al., 2022b) to solve unseen tasks after training on a large volume of instruction-tuning data, which is previously thought to be exclusive to decoder-based LLMs of large-enough sizes, can also be observed in much smaller MLMs when we do multitask Statement-Tuning with a modest amount of training data.
@@
    We explore a large number of design choices to study how Statement-Tuning benefits from the number of statement examples and statement template and task diversity in multitask Statement-Tuning, and demonstrate the data/resource-efficiency of Statement-Tuning.
@@ 2 Related Work
Zero-Shot and Few-Shot Approaches Utilizing Label Semantics
Various approaches have been proposed to reformulate zero-shot classification to leverage label semantics instead of indices, enabling the more generalized use of encoder models. In the reformulated task, textual labels are combined with the original input text, and the model should predict whether the label matches the text. TARS Halder et al. (2020) utilizes the simple concatenation of input text and label text, with no explicit connection through natural language. Similarly, Xu et al. (2023) reformulate discriminative tasks with minimal prompts, which are mostly simple concatenations of elements in the raw input. They find it effective for zero-shot generalization for ELECTRA-style encoder models but don’t test other encoder-only models with normal MLM training. Yin et al. (2019) propose an entailment-based approach where the input is a pair of texts: the original input text is the premise and label is converted into a hypothesis. However, the approach is limited by the fact that not every discriminative task can be formulated as an entailment task (in the form of a premise and a hypothesis). The entailment formulation also limits the training tasks to entailment datasets (MNLI, RTE, and FEVER), which often causes the model to over-rely on spurious lexical patterns, hindering generalization (Ma et al., 2021). Our approach Statement-Tuning serves as a more universal formulation of any discriminative task in the form of statements.
Few-shot approaches utilizing cloze-style templates have also been proposed. PET Schick and Schütze (2021a) requires an ensemble of encoder models and iterative training, relies on additional unlabeled data; Improved PET variants Schick and Schütze (2021b); Tam et al. (2021) use complex losses to compute the probability of each token in multi-token labels. Our method keeps the normal sequence classification training paradigm, while in the meantime effectively leverages the label semantics, enabling straightforward zero-shot generalization on a single encoder model.
@@
LLMs excel at unseen-task/zero-shot generalization Brown et al. (2020). Building on this, recent work explores multitask training with diverse prompts for improved zero-shot performance Sanh et al. (2022); Wei et al. (2022a); Chung et al. (2022). These methods fine-tune large models on constructed datasets with various task prompts, achieving strong zero-shot results on unseen tasks. However, effective instruction-tuning often requires billions of parameters Zhang et al. (2024), limiting their application to smaller models. Ye et al. (2022) aim to distill this zero-shot ability in a smaller model like an LSTM through synthetic data creation using an LLM, but they create task-specific models rather than a single smaller model that is capable of generalizing. Our work demonstrates similar or superior generalization than LLMs using a single smaller MLM with less training data.
@@ 3 Method: Statement-Tuning
Any discriminative task with a finite set of targets can be verbalized into a finite set of natural language statements. Figure 2 shows the example of converting the MNLI task into statements. Similar to prompting, each task has its own statement templates, based on each possible label. The truth label for training purposes on each statement depends on whether the statement contains the correct target label or not.
Statement Fine-Tuning
@@
To create the training data for multitask statement fine-tuning, we exhaustively generate statements across 16 diverse NLP datasets (categorized into 9 tasks, see Appendix E) using many varied statement templates (see Appendix A) per dataset: QQP Sharma et al. (2019), Winogrande Sakaguchi et al. (2020), PiQA Bisk et al. (2020), MNLI Williams et al. (2018), SNLI Bowman et al. (2015), Mintaka Sen et al. (2022), Yelp Polarity Zhang et al. (2015), WikiLingua Ladhak et al. (2020), SQuAD Rajpurkar et al. (2016), TweetEval’s Offensive task Zampieri et al. (2019), Massive FitzGerald et al. (2022); Bastianelli et al. (2020), Definite Pronoun Resolution Rahman and Ng (2012), QASC Khot et al. (2020), SciQ Johannes Welbl (2017), RACE Lai et al. (2017), and SAMSum Gliwa et al. (2019). We fine-tune RoBERTa Liu et al. (2019) with a binary sequence classification head to predict the truth value of the statements. By fine-tuning the model across diverse tasks, templates, and domains, the model should be able to generalize across unseen templates and tasks, as long as it can be phrased as a true/false statement.
Zero-Shot and Few-Shot Inference
@@
To perform inference on statement-finetuned RoBERTa, we also need to transform the input into statements. We randomly choose a statement template for each dataset at inference time. In our experiments, we show that our statement-tuning is robust to different templates. We exhaustively generate a statement for each possible label, as shown in Figure 1. Then, for each statement corresponding to each label, we predict the probability of such a statement being true. The final label is the statement with the highest true probability. Zero-shot inference is done by directly performing the aforementioned inference regime on the statement-finetuned RoBERTa, while K-shot inference is done after continual fine-tuning on K examples of task-specific statements.
#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion Avg
Meta-Llama-3-70B-Instruct 70B 89.0 71.3 42.0 94.7 82.7 61.9 51.8 70.5
Llama-2-13b-chat-hf 13B 89.6 60.8 40.9 93.7 82.4 53.2 51.6 67.5
Llama-2-7b-chat 7B 86.6 54.4 40.1 90.5 78.5 47.8 50.0 64.0
Mistral-7B-Instruct-v0.2 7B 89.4 73.0 41.4 88.9 82.3 57.7 55.3 69.7
Qwen1.5-7B-Chat 7B 87.0 75.5 42.1 95.3 79.7 59.1 57.8 70.9
Pythia-6.9B 6.9B 82.2 62.0 41.7 83.3 71.2 32.2 25.1 56.8
Pythia-2.8B 2.8B 79.6 68.4 41.2 77.7 69.7 12.1 35.4 54.9
Phi-2 2.7B 87.2 67.9 41.8 86.6 77.7 38.7 53.1 64.7
FlanT5-Large 770M 67.6 81.1 40.1 96.0 63.0 51.0 59.9 65.5
Qwen1.5-0.5B-Chat 500M 69.2 32.6 38.7 69.7 68.9 21.9 6.6 43.9
BART-large-mnli 406M 50.4 35.8 46.9 49.4 47.3 6.5 11.7 35.4
FlanT5-Small 60M 52.8 31.9 42.0 88.8 51.5 24.5 21.7 44.7
[…]
RoBERTa-base (Best) 125M 75.3(0.5) 72.3(1.5) 61.4(0.6) 92.9(1.3) 79.1(1.1) 40.2(3.8) 48.5(5.1) 67.1
RoBERTa-base (4k) 125M 72.4(0.5) 69.6(1.1) 60.7(0.9) 92.3(0.8) 78.5(2.7) 37.9(2.7) 46.6(4.3) 65.4
RoBERTa-large (Best) 355M 85.1(0.7) 71.8(0.8) 74.2(1.4) 95.4(0.4) 92.1(0.7) 49.9(2.1) 50.7(1.4) 75.3
RoBERTa-large (10k) 355M 85.1(0.7) 71.5(0.8) 73.0(2.4) 95.4(0.4) 91.1(0.8) 48.4(0.7) 49.1(3.2) 73.4
Full/3000-shot:
RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2 -
RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2 -
Table 1: Comparison of our approach against many pre-trained open-source encoder-decoder and decoder-only Large Language Models on 7 Natural Language Understanding tasks in zero-shot conditions. FT stands for Full Finetuning and is included for reference. For Statement-Tuning, we report the average across 5 training runs and 5 evaluation runs and include the average standard deviation in parenthesis. We highlight all scores in gray where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
@@ 4 Experimental Setup
4.1 Evaluation Datasets
We measure our model’s generalizability using another set of 7 diverse datasets representing a variety of unseen tasks or unseen domains: Balanced COPA (BCOPA; Kavumba et al., 2019; Roemmele et al., 2011), MRPC Dolan and Brockett (2005), Emotion Saravia et al. (2018), Amazon Polarity McAuley and Leskovec (2013); Zhang et al. (2015), FigQA Liu et al. (2022), StoryCloze (2016) Yang et al. (2023), and Yahoo Answers Topics Zhang et al. (2015). Among the evaluation data, MRPC (paraphrase identification) and Amazon Polarity (sentiment analysis) represent tasks seen during training but in different domains and demonstrate cross-domain generalizability. The rest are unseen tasks and hence examine the cross-task generalizability.
@@ 4.2 Statement Finetuning Configurations
We statement-finetune both RoBERTa-base and RoBERTa-large across diverse NLP tasks outlined in Section 3. However, as statement fine-tuning expands the dataset with various templates over all possible labels, it is arguably unwise to fine-tune on all possible generated statements. Moreover, each task has a different data size, leading to unbalanced fine-tuning data. Therefore, we sample statements randomly for each task, uniformly across true and false statements. In true/false statements, we also balance original classes. We explore sample size from 1,000 statements to 50,000 statements per dataset. We encourage the invariance to phrasing in diverse statements by designing multiple statement templates per dataset (a list of all statement templates is shown in Appendix A). Furthermore, we run the training five times to account for randomness in training data creation. In the evaluation, we randomly pick a template for each dataset in a single evaluation run and also repeat the evaluation five times. We thus report the mean and standard deviation of 5 $\times$× 5 runs to show the general task accuracy and the (in)variance to phrasing. We also explore the effect of statement diversity during training in Section 5.2.3.
After multitask statement tuning is completed, we can further continue fine-tuning the model on the target downstream dataset. Specifically, we explore various n-shot configurations: Full/3,000-shot, 1,000-shot, 500-shot, 200-shot, and 32-shot, where we use limited data from the training sets of the corresponding dataset to fine-tune our statement-tuned models. For the Full/3,000-shot case, we cap the training set at 3,000 examples, otherwise, we use the entire set (this is the case for Amazon Polarity only). For StoryCloze, there is no training set, so we just carry out 32-shot (using 32 samples from the test set for fine-tuning and evaluating on the rest) and zero-shot experiments. As for Yahoo Answers Topic and Emotion, due to them being multi-class classification tasks, we cap the n-shot analysis at 200-shot due to the larger number of choices per example (and hence a larger number of statements per example).